# Enhancing noise characterization with robust time delay interferometry combination

### v2

@@ Abstract
Time delay interferometry (TDI) is essential for suppressing laser frequency noise and achieving the targeted sensitivity for space-borne gravitational wave (GW) missions. In Paper I, we examined the performance of the fiducial second-generation TDI Michelson configuration versus an alternative, the hybrid Relay, in noise suppression and data analysis. The results showed that both TDI schemes have comparable performances in mitigating laser and clock noises. However, when analyzing chirp signal from the coalescence of massive binary black holes, the Michelson configuration becomes inferior due to its vulnerable T channel and numerous null frequencies. In contrast, the hybrid Relay is more robust in dynamic unequal-arm scenarios. In this work, we further investigate the noise characterization capabilities of these two TDI configurations. Our investigations demonstrate that hybrid Relay achieves more robust noise parameter inference than the Michelson configuration. Moreover, the performance could be enhanced by replacing the T channel of hybrid Relay with the null stream from Sagnac configuration. The combined data streams, two science observables from the hybrid Relay and a null observable from the Sagnac, could form an optimal dataset for characterizing noises.
@@ I Introduction
In this work, we further investigate the capability of hybrid Relay in noise characterization and compare it with the fiducial Michelson. The results show that the T channel is crucial for breaking the degeneracy between optical metrology system (OMS) noises in the science channels and for precisely determining noise parameters. The performance of Michelson configuration is still limited by its unstable T channel and numerous null frequencies, whereas the hybrid Relay shows potential for superior capability. However, the performance of hybrid Relay is also affected by its T observable, which is subject to increased null frequencies. To enhance its capability, the T from the Sagnac configuration is selected to replace the hybrid Relay-T, enabling noise characterization with hybrid Relay-A and E channels. This combination of data streams could enhance characterization efficiency.
@@ II Time delay interferometry
For the optimal data streams, A and E can effectively respond to GW signals, with their antenna patterns equivalent to two orthogonal interferometer rotated by [15]. The T channel will be a noise-dominated data stream, especially at frequencies lower than 50 mHz. As detailed in the following section, the T channel plays a crucial role in noise characterization.
@@ III Correlations between TDI and spectra stability
The second-generation TDI observables can effectively suppress laser frequency noises [26, 24, 27, 10, 28, 29, and references therein] and mitigate clock noise [30]. Acceleration noise and OMS noise are expected to be dominating noises in the data streams, and their respective noise spectra could be [1],
where represents the amplitude of acceleration noise, and represents the amplitude of OMS noise. Assuming identical noise characteristics, the amplitudes are set to be and for all measurement system on each S/C.
@@
Conversely, for hybrid Relay, three data streams demonstrate greater independence from each other, as indicated by the dashed lines in lower plots. The highest correlation between E and T is less than 0.03 in the frequency band below 10 mHz. T exhibits high correlation with A/E at its characteristic frequencies, which can be as low as 15 mHz. The correlations between Tα1 and hybrid Relay observables are depicted by solid lines in the lower two plots. As shown by the green lines, these two T data steams are highly correlated except at few particular frequencies. Tα1 has low relevance with A/E until 40 mHz, with correlation only increasing at its null frequencies Hz. The correlations could vary if the noises are non-identical, we evaluated such a case in Appendix A.
@@
We randomly selected one month to evaluate the PSDs of (A, E, T) channels at different time points for Michelson, Sagnac, and hybrid Relay. The ratios between the spectra and the averaged spectra are calculated, and the results are shown in Fig. 4. The upper two plots depict the deviations of two science channels for three TDI configurations, with their PSDs deviating from the average spectra around their respective null frequencies. Compared to Michelson observables, which are subject to the null frequencies at Hz, the hybrid Relay only have null frequencies at Hz, and Sagnac observables have null frequencies at Hz. Therefore, two science channels of hybrid Relay are the most robust. The ratios for three T channels are shown in the lower panel of Fig. 4. For the TX1 observable from Michelson, besides its PSD being unstable around its characteristic frequencies, its spectrum also exhibits large deviation at frequencies lower than 2 mHz. The T channel of hybrid Relay has more null frequencies than its A/E channel with the lowest being Hz. And the Tα1 has the fewest null frequencies among three null data streams.
An optimal data stream combination would maximize the frequency band with robust spectra. Additionally, the levels of correlation between optimal channels will not affect the analysis result if a covariance matrix is employed to calculate the likelihood, as shown in Eqs. (14)-(15). Lower correlations between observables would reduce the error of signal-to-noise ratio calculation when the off-diagonal terms are ignored. Based on evaluations of spectral robustness and channel correlation, the data streams, (A, E, Tα1) could be an optimal combination for the data analysis. In the next section, noise characterization comparisons will be performed for different data combinations.
@@ IV Noise characterizations
During the characterization, two adjustable factors are considered: the effective frequency band and the duration of the noise data. Regarding the frequency band, the low-frequency cutoff is set to be 0.03 mHz, and the high frequency cutoff varies depending on the specific evaluations. The frequency range of [0.03, 10] mHz is considered a ’clean’ band, where the correlation between three optimal channels are steady and the spectra remain stable over a month’s duration (except the Michelson-T channel), as shown in Figs. 3 and 4. The duration of the data is also plays a crucial role in determining the amount of data and the frequency resolution. Opting for a long duration increases spectral resolution but may introduce more fluctuations. Therefore, a trade-off is necessary when selecting the data duration and frequency range. Table 2 provides a checklist for parameter inference with various data setups and the feasibility. The inferred values of and ( represents the noise component on S/C facing S/C) are selected to represent uncertainties of the parameters. Checkmarks and crossmarks indicate whether the inference correctly or incorrectly encompass the input true values.
@@
Based on these analysis, the T channel proves to be the ’shortest plank in the barrel’ in both Michelson and hybrid Relay. In Michelson configuration, the PSD/CSD involving TX1 channel fluctuates significantly with mission time in the lower frequency band, undermining Michelson’s capability for noise characterizing. On the other hand, in the hybrid Relay configuration, T exhibits more unstable null frequencies than its science channels as illustrated in Fig. 4, and its correlations with A and E soar around these null frequencies. Tα1 from Sagnac exhibits more robust noise spectra compared to the T, making it a better substitute for T given their high correlation. In this case, extending the high frequency cutoff to 30 mHz avoids unstable null frequency for three optimal observables . The inference results from this combination using 120 days and 180 days data with a 0.03 Hz high frequency cutoff are shown in Fig. 8. The inferred distributions from 120 days accurately and precisely encompass the true values compared to previous data setups, with uncertainties of the parameters and the corresponding spectra less than a few percents of the true values. The inference precisions from 180 days data further increase, but with a few parameters outside the credible regions.
@@ V Conclusion and discussion
In this work, we compare the noise characterizations by varying frequency bands and data duration with different data combinations for Michelson and hybrid Relay TDI configurations. Our findings highlight that the hybrid Relay is more effective in determining noise parameters compared to the Michelson configuration. The performance of Michelson is constrained by the unstable spectra with T channel and fluctuations around numerous null frequencies during the orbital evolution of detector. In contrast, the noise spectra of science observables from hybrid Relay are more robust across a wider frequency range due to the reduced occurrence of null frequencies. The efficiency of hybrid Relay can be further enhanced by replacing its T observable with the null channel from Sagnac configuration, where two observables are highly correlated and the latter has fewer null frequencies than the former.
In paper I, we evaluated the performance of two TDI configurations in analyzing chirp signals from massive black hole binary coalescences, revealing deficiencies in the Michelson TDI observables. In that context, the T channel plays a negligible role as the analysis was primarily influenced by the two science channel, A and E. Our current investigation underscores the critical importance of a reliable null channel in breaking degeneracies between noises and accurately determining noise parameters. To enhance this, we have opted to integrate the T channel from Sagnac with the A and E from hybrid Relay, forming an optimal dataset, . This selection is made conservatively in the second-generation TDI observables, considering the sufficient suppression of laser noise and clock noise in TDI channels. Given the diversity and abundance of TDI observables, there may exist even more robust null observable suitable for noise characterization with ().
@@ Acknowledgements.
GW was supported by the National Key R&D Program of China under Grant No. 2021YFC2201903, and NSFC No. 12003059. This work made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory. This work are performed by using the python packages [38], [39], [40], [36] and [37], and the plots are make by utilizing [41], [42].Appendix A Correlations between TDI with non-identical noise setup

### v3

@@ Abstract
Time delay interferometry (TDI) is essential for suppressing laser frequency noise and achieving the targeted sensitivity for space-borne gravitational wave (GW) missions. In Paper I, we examined the performance of the fiducial second-generation TDI Michelson configuration versus an alternative, the hybrid Relay, in noise suppression and data analysis. The results showed that both TDI schemes have comparable performances in mitigating laser and clock noises. However, when analyzing chirp signal from the coalescence of massive binary black holes, the Michelson configuration becomes inferior due to its vulnerable T channel and numerous null frequencies. In contrast, the hybrid Relay is more robust in dynamic unequal-arm scenarios. In this work, we further investigate the noise characterization capabilities of these two TDI configurations. Our investigations demonstrate that hybrid Relay achieves more robust noise parameter inference than the Michelson configuration. Moreover, the performance can be enhanced by replacing the T channel of hybrid Relay with a second-generation TDI null stream . The combined three data streams, including two science observables from the hybrid Relay and , could reduce the instabilities of noise spectra in the targeting frequency band and form an optimal dataset for characterizing noises.
@@ I Introduction
In this work, we further investigate the capability of hybrid Relay in noise characterization and compare it with the fiducial Michelson. The results show that the T channel is crucial for breaking the degeneracy between optical metrology system (OMS) noises in the science channels and for precisely determining noise parameters. The performance of Michelson configuration is still limited by its unstable T channel and numerous null frequencies, whereas the hybrid Relay shows potential for superior capability. However, the performance of hybrid Relay is also affected by its T observable, which is subject to increased null frequencies. To enhance its capability, the null stream identified in [21] is selected to replace the hybrid Relay-T. The combined three data streams, A and E of the hybrid Relay and , can reduce the instability of noise spectra in the targeting frequency band and improve efficiency of noise characterization.
@@ II Time delay interferometry
![Refer to caption](x3.png)
For the optimal data streams, A and E can effectively respond to GW signals, with their antenna patterns equivalent to two orthogonal interferometer rotated by [15]. The T channel will be a noise-dominated data stream, especially at frequencies lower than 50 mHz. As detailed in the following section, the T channel plays a crucial role in noise characterization as null stream.
Besides the T channels by combining three ordinary channels, the specific null streams have been developed. Typical cases include the first-generation TDI fully symmetric Sagnac [4] and the the second-generation TDI [6]. In the unequal arms case, replacing T channel in the first-generation TDI Michelson with could enhance noise characterization and stochastic GW reconstruction [14]. Hartwig and Muratore [21] developed more second-generation TDI null streams and and examined their noise spectra. Among these null streams, has null frequencies at , and has no null frequency in target frequency band. Therefore, these two null observables are selected as potential substitutes for the T channels. The geometries of the are shown in Fig. 3.
![Refer to caption](x4.png)
@@ III Correlations between TDI and spectra stability
![Refer to caption](x5.png)
![Refer to caption](x6.png)
![Refer to caption](x7.png)
![Refer to caption](x8.png)
@@
The second-generation TDI observables can effectively suppress laser frequency noises [27, 24, 28, 10, 29, 30, and references therein] and mitigate clock noise [31]. Acceleration noise and OMS noise are expected to be dominating noises in the data streams, and their respective noise spectra could be [1],
where and represent the amplitudes of acceleration noise and OMS noise, respectively. Assuming identical noise characteristics, the amplitudes are set to be and for all measurement system on each S/C.
@@
Conversely, for hybrid Relay, three data streams demonstrate greater independence from each other, as indicated by the dashed lines in lower plots. The highest correlation between E and T is less than 0.03 in the frequency band below 10 mHz. T exhibits high correlation with A/E at its characteristic frequencies, which can be as low as 15 mHz. The correlations between and hybrid Relay observables are depicted by solid lines in the lower two plots. As shown by the green lines, these two data steams are highly correlated except at a few particular frequencies. has low relevance with A/E, and their correlations only increase at frequencies Hz. The correlations could vary if the noises are non-identical, we evaluated such a case in Appendix A.
![Refer to caption](x9.png)
![Refer to caption](x10.png)
![Refer to caption](x11.png)
![Refer to caption](x12.png)
@@
We randomly selected one month to evaluate the PSDs of (A, E, T) channels at different time points for Michelson, Sagnac, and hybrid Relay, as well as the null streams and . The ratios between the spectra and the averaged spectra are calculated, and the results are shown in Fig. 5. The upper two plots depict the deviations of two science channels for three TDI configurations, with their PSDs deviating from the average spectra around their respective null frequencies. Compared to Michelson observables, which are subject to the null frequencies at Hz, the hybrid Relay only have null frequencies at Hz, and Sagnac observables have null frequencies at Hz. Therefore, two science channels of hybrid Relay are the most robust. The ratios for three T channels are shown in the lower left panel of Fig. 5. For the TX1 observable from Michelson, besides its PSD being unstable around its characteristic frequencies, its spectrum also exhibits instability at frequencies lower than 2 mHz. The T channel of hybrid Relay has more null frequencies than its A/E channel with the lowest being Hz, and the Tα1 has the unstable noise spectrum around its null frequencies. The PSD ratio of is change with time-varying arm lengths in all targeting band, while the only has fluctuation around the and is the most stable one among five null data streams. The noise PSD of are formulated and plotted in Appendix C.
An optimal data stream combination would maximize the frequency band with robust spectra. Additionally, the levels of correlation between optimal channels will not affect the analysis result if a covariance matrix is employed to calculate the likelihood, as shown in Eqs. (14)-(15). Lower correlations between observables would reduce the error of signal-to-noise ratio calculation when the off-diagonal terms are ignored. Based on evaluations of spectral robustness and channel correlation, the data streams, (A, E, ) could be an optimal combination for the data analysis as shown in the lower right plot of Fig.5. In the next section, noise characterization comparisons will be performed for different data combinations.
@@ IV Noise characterizations
During the characterization, two adjustable factors are considered: the effective frequency band and the duration of the noise data. Regarding the frequency band, the low-frequency cutoff is set to be 0.03 mHz, and the high frequency cutoff varies depending on the specific evaluations. The frequency range of [0.03, 10] mHz is considered a ’clean’ band, where the correlation between three optimal channels are steady and the spectra remain stable over a month’s duration (except the Michelson-T channel), as shown in Figs. 4 and 5. The duration of the data is also plays a crucial role in determining the amount of data and the frequency resolution. Opting for a long duration increases spectral resolution but may introduce more fluctuations. Therefore, a trade-off is necessary when selecting the data duration and frequency range. Table 2 provides a checklist for parameter inference with various data setups and the feasibility. The inferred values of and ( represents the noise component on S/C facing S/C) are selected to represent uncertainties of the parameters. Checkmarks and crossmarks indicate whether the inference correctly or incorrectly encompass the input true values.
@@
![Refer to caption](x13.png)
![Refer to caption](x14.png)
![Refer to caption](x15.png)
![Refer to caption](x16.png)
@@
![Refer to caption](x17.png)
![Refer to caption](x18.png)
![Refer to caption](x19.png)
![Refer to caption](x20.png)
Based on these analysis, the T channel proves to be the ’shortest plank in the barrel’ in both Michelson and hybrid Relay. In Michelson configuration, the PSD/CSD involving TX1 channel fluctuates significantly with mission time in the lower frequency band, undermining Michelson’s capability for noise characterizing. On the other hand, in the hybrid Relay configuration, T exhibits more unstable null frequencies than its science channels as illustrated in Fig. 5, and its correlations with A and E soar around these null frequencies. Tα1 from Sagnac exhibits more robust noise spectra compared to the T, making it a substitute for T given their high correlation. In this case, extending the high frequency cutoff to 30 mHz avoids unstable null frequency for three optimal observables . The inference results from this combination using 120 days with a 0.03 Hz high frequency cutoff can yield sensible parameter distributions, but using 180 days of data fails to accurately infer all 12 parameters.
As the optimal combination, the results of dataset using 180 days with 0.1 Hz high frequency cutoff are shown in Fig. 9. The inferred distributions accurately and precisely encompass the true values compared to previous data combinations, with uncertainties being less than a percents of the true values for OMS noise. Compare to the result from three optimal TDI observables from hybrid Relay with 180-day data, as shown in Table 2, the novel combination significantly reduce the uncertainties of the OMS noise parameters by 10 times, benefiting from the extended high frequency cutoff from 0.01 Hz to 0.1 Hz. This improvement is much greater compared to the correctly inferred results from the two Michelson science channels, (AX1, EX1). The constraints on the parameters of acceleration noise are not improved since the OMS noise dominates frequency band [0.01, 0.1] Hz.
@@ V Conclusion and discussion
In this work, we compare the noise characterizations by varying frequency bands and data duration with different data combinations for Michelson and hybrid Relay TDI configurations. Our findings highlight that the hybrid Relay is more effective in determining noise parameters compared to the Michelson configuration. The performance of Michelson is constrained by the unstable spectra with T channel and fluctuations around numerous null frequencies during the orbital evolution of detector. In contrast, the noise spectra of science observables from hybrid Relay are more robust across a wider frequency range due to the reduced occurrence of null frequencies. The efficiency of hybrid Relay can be further enhanced by replacing its T observable with the null stream , where two observables are highly correlated and the latter has as few null frequencies as the science observables of hybrid Relay.
In paper I, we evaluated the performance of two TDI configurations in analyzing chirp signals from massive black hole binary coalescences, revealing deficiencies in the Michelson TDI observables. In that context, the T or null channel plays a negligible role as the analysis was primarily influenced by the two science channel A and E. Our current investigation underscores the critical importance of a reliable null channel in breaking degeneracies between noises and accurately determining noise parameters.
@@ Acknowledgements.
G.W. acknowledges Olaf Hartwig’s valuable comments, which resulted in the the inclusion of the null stream into the optimal dataset. G.W. also thanks Martina Muratore, Quentin Baghi, Jean-Baptiste Bayle, Shichao Wu for their helpful discussions during the 15th International LISA Symposium 2024. G.W. was supported by the National Key R&D Program of China under Grant No. 2021YFC2201903, and NSFC No. 12003059. This work made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory. This work are performed by using the python packages [39], [40], [41], [37] and [38], and the plots are make by utilizing [42], [43].Appendix A Correlations between TDI with non-identical noise setup
@@
![Refer to caption](x23.png)
![Refer to caption](x24.png)
Appendix C PSD of null stream
The noise PSD of , including acceleration noise and OMS noise, is expressed as
where is the arm length. As depicted in Fig. 12, its noise spectrum only has the same null frequencies as the science channel A and E of hybrid Relay.
![Refer to caption](x25.png)